### v3

# RankCLIP: Ranking-Consistent Language-Image Pretraining

Abstract
Self-supervised contrastive learning models, such as CLIP, have set new benchmarks for vision-language models in many downstream tasks. However, their dependency on rigid one-to-one mappings overlooks the complex and often multifaceted relationships between and within texts and images. To this end, we introduce RankCLIP, a novel pretraining method that extends beyond the rigid one-to-one matching framework of CLIP and its variants. By extending the traditional pair-wise loss to list-wise, and leveraging both in-modal and cross-modal ranking consistency, RankCLIP improves the alignment process, enabling it to capture the nuanced many-to-many relationships between and within each modality. Through comprehensive experiments, we demonstrate the effectiveness of RankCLIP in various downstream tasks, notably achieving significant gains in zero-shot classifications over state-of-the-art methods, underscoring the importance of this enhanced learning process. Code and model checkpoints are available at https://github.com/Jam1ezhang/RankCLIP.
1 Introduction
In the realm of computer vision (CV) [55], natural language processing (NLP) [7], and multimodal deep learning [25, 60, 4], the alignment between visual and textual modalities [49, 5] has emerged as a cornerstone for downstream applications, ranging from image captioning [18] to zero-shot classification [42].
Contrastive Language-Image Pretraining (CLIP) [43] marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand [22, 23, 6] and generate [44, 11] descriptions of visual contents. Following the success of this contrastive learning paradigm, many recent works have been developed upon the original CLIP. More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision [28], as well as improving downstream performance via cross-modal late interaction [59], hierarchical feature alignment [15], geometric consistency regularization [19], additional learning [36], adaptive loss [58], hierarchy-aware attentions [17], and softer cross-modal alignment [16].
Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data [8]. For example, as shown in Fig. 1, while pretrained models like CLIP can correctly classify dog, cat and airplane, they do not necessarily learn that dog and cat are more close to each other than dog and airplane, in terms of both in-modal (dog text is more similar to cat text than to airplane text) and cross-modal (dog text is more matched to cat image than to airplane image) similarities. Since it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, a large amount of information not used and unknown to the model during and after the training process.
Recognizing the complex many-to-many relationships as well as the rich information contained within both in-modal and cross-modal data, we introduce Ranking-Consistent Language-Image Pretraining, (RankCLIP), which employs ranking consistency to learn and optimize similarity levels both between (cross-modal) and within (in-modal) the text-image pairs. The concept of ranking consistency stems from the simple observations that similar texts often correlate with similar images, as seen with the dog, cat and airplane example in Fig. 1. It effectively captures secondary similarity relationships among unmatched pairs, enabling the model to learn more efficiently for free compared to relying solely on matched pairs. Ranking consistency is conveniently modeled as an additional loss term to the traditional contrastive loss, requiring no extra external modules. It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency [28], potentially boosting performance in both efficiency and effectiveness.
The main contributions of this paper are: 1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and 2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
2 Related Work
Vision-language pretraining has witnessed significant advancements over the past years [3, 14, 31]. Models such as CLIP [43], ALIGN [26] and FLAVA [49] have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks. Many follow-up works, mostly built upon CLIP, have been proposed since then. Li et al. [28] introduced DeCLIP, improving zero-shot performance through intrinsic supervision. FILIP [59] advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism. Gao et al. [15] developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance. Additionally, SLIP [36] merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy. Goel et al. [19] introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
Recently, Yang et al. [58] introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments. HiCLIP [17] refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts. EqSim [57] incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs. Additionally, SoftCLIP [16] softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
Compared with existing approaches, RankCLIP sets itself apart by fully leveraging the many-to-many relationships within each batch of text-image pairs, promoting learning from both matched and unmatched pairs with varying similarities by integrating in-modal and cross-modal list-wise ranking consistencies into the contrastive training objective. Crucially, RankCLIP diverges from existing models’ pair-wise training objective by adopting a global, list-wise optimization approach. In other words, it considers the rankings of all images and texts collectively within each batch, rather than focusing on pairwise similarities as seen in other methods.
3 RankCLIP
RankCLIP efficiently leverages the many-to-many relationships in real-world data by focusing on both matched and unmatched pairs. As in Fig. 2, it not only identifies if an image-text pair matches but also assesses their relative semantic similarities to other images and texts of both modalities in the dataset through self-supervised ranking consistency. Uniquely, RankCLIP employs a list-wise loss for training batches, distinguishing it from other methods that solely rely on pair-wise relationships, as discussed in §2.
3.1 Ranking Model Formulation
RankCLIP leverages the Plackett-Luce (PL) ranking model [41, 32, 20] to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured. Specifically, for a given data pair, whether it is in-modal (image-image, text-text), or cross-modal (image-text), we calculate its in- or cross-modal cosine similarity to serve as the score when measuring the alignment of its ranking with respect to another reference ranking .
Following [41], we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from . The probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
where denotes the set of items ranked before . Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
RankCLIP’s objective is to maximize the consistency log-likelihood of the list ranking in one modality towards the reference ranking (from the same/in-modal and different/cross-modal data), which conveniently aligns with minimizing the negative log-likelihood loss:
3.2 Cross-modal Consistency Ranking
As illustrated by the green box in Fig. 2, RankCLIP utilizes secondary relationships between unmatched visual and textual representations by constructing a list-wise rank loss. This approach ensures that the semantic similarity rankings between one image and multiple texts align with those between one corresponding text and multiple images. For example, as shown in Fig. 1, from the dog perspective, the semantic distance between dog image and cat text is closer compared to the plane text. This relationship should also apply between the dog text and the cat, plane images. Mathematically, Eq. (3) can be specified as:
By optimizing Eq. (4), RankCLIP enhances its ability to bridge the semantic gap between modalities by leveraging nuanced unmatched correlations. This can also be viewed as learning a symmetric cosine-similarity matrix, further reinforcing semantic consistency across modalities.
ImageNet1K | MSCOCO | ||||||||
Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||
Top-1 | Top-3 | Top-5 | Recall@1 | Recall@5 | Recall@10 | Recall@1 | Recall@5 | Recall@10 | |
CLIP [43] | 9.06% | 16.94% | 21.63% | 6.68% | 18.36% | 26.94% | 3.70% | 9.74% | 14.04% |
CyCLIP [19] | 9.40% | 17.32% | 21.72% | 6.50% | 19.34% | 29.14% | 3.72% | 11.16% | 16.06% |
ALIP [58] | 9.71% | 18.31% | 23.07% | 6.04% | 18.04% | 26.92% | 3.70% | 10.22% | 14.38% |
RankCLIP | 10.16% | 19.57% | 24.01% | 7.18% | 19.46% | 28.48% | 3.74% | 10.28% | 14.18% |
3.3 In-modal Consistency Ranking
The pink box in Fig. 2 highlights the in-modal consistency component of the proposed rank loss. RankCLIP ensures semantic consistency within each modality – image to image and text to text – enhancing the use of secondary unmatched relationships as an optimization objective. The underlying principle is that similar images should correspond to similar texts. For example, in Fig. 1, from the dog image perspective, the cat image is the most similar, followed by the plane image. This relationship should hold true for their corresponding texts as well, where we utilize this to construct our and from Eq. (3). Mathematically, Eq. (3) can be specified as:
where and are the text and image batch embedding matrix, respectively. Via Eq. (6), the model can efficiently leverage the nuanced in-modal relationships to learn a richer and more structured semantic representation.
3.4 RankCLIP Loss
Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix B), the complete rank loss is thus formulated as:
which is also depicted in Fig. 2. By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
3.5 Training Recipe on Selecting and
In the early stage of pre-training, rank consistency is highly unstable due to random initialization. Overemphasizing ranking consistency at this stage can impede the optimization of the embedding space. To address this, we gradually increase the weights and of the ranking loss as training progresses. Specifically, we have:
where and denote the current training epoch and total number of epoch, respectively. The full RankCLIP framework is outlined in Algorithm 1.
4 Experiments
4.1 Experimental Setup
Baselines. The most direct baseline to RankCLIP is the original CLIP [43]. To further demonstrate the superior performance of RankCLIP, we include CyCLIP [19], which introduces cyclic consistency constraints to enforce more robust alignment between visual and textual representations, improving generalization and semantic coherence. We also include ALIP [58], which leverages synthetic captions to enhance vision-language representation learning. More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, making its approach complementary to ours. The training procedures and parameters for all models are detailed in Appendix A.
Data. All approaches are pretrained on the Conceptual Captions 3M (CC3M) dataset [48], which contains approximately 3.3 million text-image pairs. Although significantly smaller than CLIP’s original dataset [24], CC3M remains a standard benchmark for vision-language pretraining, enabling strong zero-shot performance [2, 28, 53, 36, 19]. As discussed in §5.2, we also train CLIP and RankCLIP on a larger 15M-image subset of YFCC100M [54] (YFCC15M) to study the effect of scaling up the dataset size.
ImageNetV2-Matched | ImageNetV2-Threshold | ImageNetV2-Top | ImageNet-R | |||||||||
Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
CLIP [43] | 7.53% | 14.99% | 19.61% | 8.89% | 17.22% | 21.86% | 10.76% | 19.80% | 24.87% | 9.36% | 10.56% | 19.76% |
CyCLIP [19] | 7.68% | 15.07% | 19.11% | 9.10% | 17.42% | 21.94% | 11.20% | 20.18% | 25.34% | 9.23% | 16.72% | 21.64% |
ALIP [58] | 7.82% | 15.56% | 19.81% | 9.65% | 18.31% | 22.85% | 11.43% | 20.88% | 26.10% | 10.92% | 20.27% | 26.24% |
RankCLIP | 9.01% | 16.95% | 21.12% | 10.32% | 19.31% | 24.13% | 12.31% | 22.11% | 27.17% | 11.34% | 20.88% | 26.94% |
CIFAR-10 |
CIFAR-100 |
DTD |
FGVGAircraft |
Food101 |
GTSRB |
OxfordPets |
SST2 |
STL10 |
SVHN |
Average |
CLIP [43] | 77.6% | 56.2% | 43.2% | 22.6% | 39.7% | 60.0% | 40.4% | 51.0% | 79.0% | 50.5% | 52.0% |
CyCLIP [19] | 76.8% | 54.3% | 45.8% | 19.2% | 37.5% | 58.6% | 44.2% | 51.5% | 82.3% | 41.3% | 51.2% |
ALIP [58] | 71.1% | 49.1% | 47.1% | 17.4% | 36.1% | 51.5% | 41.9% | 53.3% | 81.0% | 38.3% | 48.7% |
RankCLIP | 78.4% | 56.6% | 42.4% | 23.4% | 40.2% | 60.6% | 40.6% | 53.4% | 79.6% | 47.7% | 52.3% |
4.2 Zero-shot Classification
We evaluate the zero-shot classification performance of CLIP [43], CyCLIP [19], ALIP [58], and RankCLIP on ImageNet1K [12, 46]. As shown in Table 1, RankCLIP consistently outperforms CLIP, which highlight the effectiveness of ranking consistency in enhancing language-image alignment with the same training data. Compared to CyCLIP [19], which enforces cyclic consistency to improve semantic coherence, RankCLIP achieves higher accuracy across all metrics, suggesting that ranking consistency provides a more direct and effective regularization for representation learning. Additionally, RankCLIP surpasses ALIP [58], indicating that ranking consistency is a stronger alternative to synthetic caption-based supervision. Notably, RankCLIP shows the most significant improvement in top-1 accuracy, reinforcing its practical advantages where the highest-ranked prediction is most critical.
4.3 Zero-shot Cross-modal Retrieval
We further evaluate RankCLIP on zero-shot cross-modal retrieval tasks, including image-to-text and text-to-image retrieval, using the MSCOCO [30] dataset. As shown in Table 1, RankCLIP outperforms all baselines, though the improvements are less pronounced compared to zero-shot classification. This smaller margin of improvement may be attributed to the increased complexity of retrieval tasks, which require fine-grained image-text alignment across varying resolutions and object details – challenges distinct from the more direct pattern recognition in classification.
4.4 Robustness to Distribution Shifts
To evaluate the robustness of RankCLIP under distribution shifts, we test all approaches on three variants of ImageNetV2 [45] and ImageNet-R [22], which assess resilience to different real-world deviations from ImageNet1K. As shown in Table 2, RankCLIP consistently outperforms all baselines, achieving the highest accuracy across all datasets. These results indicate that RankCLIP not only improves standard zero-shot classification but also enhances adaptation to real-world distribution shifts.
4.5 Linear Probing
We further assess whether the advantages of ranking consistency persist when supplemented with in-domain supervision. Specifically, we apply linear probing, where pretrained encoders remain fixed while a logistic regression classifier is trained on domain-specific datasets. We evaluate on 10 standard image classification benchmarks, including CIFAR-10, CIFAR-100, DTD [9], FGVG-Aircraft [33], Food101 [1], GTSDB [52], OxfordPets [40], SST2 [50], STL-10 [10], and SVHN [37]. As shown in Table 3, RankCLIP achieves the highest average accuracy, demonstrating that ranking consistency enhances generalization even with additional in-domain supervision.
5 Ablation Studies
5.1 Different Weights of RankCLIP Loss
In Eq. (8), we define the RankCLIP loss as a linear combination of the original contrastive loss and the in-modal and cross-modal ranking consistency losses, weighted by and , respectively. In §3.5, we introduce a training strategy that adaptively adjusts and at different stages of training. In this section, we analyze the impact of these weights and demonstrates that adaptive weighting further enhances RankCLIP ’s performance. Notice that all model variants follow the same pretraining setup detailed in Appendix A. As shown in Fig. 3, RankCLIP outperforms CLIP even with fixed and , highlighting the effectiveness of ranking consistency. And the adaptive weighting strategy further boosts accuracy by preventing the under-utilization of ranking consistency at low weights and avoiding disruptions to contrastive learning at high weights.
Method | Vision Backbone | ImageNet1K | MSCOCO | Linear Probing Avg. Acc. | |||||||
Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||||
Top-1 | Top-3 | Top-5 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |||
CLIP [43] | RN50 | 21.6% | 36.9% | 44.9% | 15.6% | 36.4% | 48.4% | 6.7% | 15.2% | 20.1% | 64.2% |
RankCLIP | 30.9% | 49.4% | 57.6% | 19.5% | 42.6% | 54.8% | 7.5% | 16.2% | 21.6% | 68.9% | |
CLIP [43] | ViT-B/32 | 20.7% | 35.0% | 42.4% | 11.9% | 29.4% | 40.8% | 5.1% | 12.9% | 17.9% | 60.7% |
RankCLIP | 26.2% | 41.4% | 48.9% | 13.8% | 33.8% | 45.9% | 6.0% | 13.6% | 18.6% | 61.3% |
5.2 Different Data Sizes
To assess the scalability of RankCLIP, we trained both CLIP and RankCLIP on 500k, 750k, 1M, and 3M text-image pairs from CC3M, as well as 15M pairs from YFCC15M, following the procedure in Appendix A. Fig. 4 compares their performance on zero-shot top-1 classification accuracy for ImageNet1K (left) and averaged linear probing results (middle), where RankCLIP consistently outperforms CLIP. Full, non-averaged linear probing results are provided in Appendix 7. Notably, the performance gains of RankCLIP become more pronounced as dataset size scales from 1m to 15m, highlighting its superior scalability, which is critical for language-image pretraining.
Fig. 4 (right) further illustrates RankCLIP’s robustness across different dataset sizes. The horizontal axis represents top-1 accuracy on standard ImageNet1K, while the vertical axis shows accuracy on ImageNet1K-R. The black diagonal () denotes ideal robustness, with deviations below it indicating degradation under distribution shifts. RankCLIP remains well above the red baseline, which represents typical in-distribution to out-of-distribution generalization [35], and stays close to the ideal line, demonstrating strong robustness.
5.3 Different Backbones: RN50 vs. ViT
We compare RankCLIP and CLIP across ResNet-50 (RN50) and ViT-B/32 backbones to assess its generalization. As shown in Table 4, RankCLIP consistently outperforms CLIP across all tasks. Specifically, for zero-shot classification on ImageNet1K, RankCLIP improves top-1 accuracy by +9.3% with RN50 and +5.5% with ViT-B/32, demonstrating stronger feature learning. In cross-modal retrieval, RankCLIP achieves +3.9% in image-to-text and +0.8% in text-to-image with RN50, with smaller but consistent gains for ViT-B/32. RankCLIP also improves linear probing accuracy by +4.7% with RN50 and +0.6% with ViT-B/32, confirming its advantage in representation learning. While both architectures benefit, RN50 sees the largest gains, suggesting that ranking consistency particularly enhances hierarchical feature extraction in CNNs.
6 Analysis
6.1 Modality Gap
We analyze the modality gaps of CLIP and RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP [34], and presenting a histogram of the gaps. The modality gap [29] refers to the separation between text and image embeddings in multimodal models, hindering joint representation learning. This gap, inherent from initialization and reinforced by contrastive learning in CLIP, challenges effective language-image modeling. Recent studies [51, 27, 38] suggest that reducing this gap improves multimodal representations and downstream performance. As shown in Fig. 6, RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances text-image alignment.
6.2 Alignment and Uniformity
Beyond reducing modality gap, effective contrastive learning should ensure a broad and uniform distribution over a hypersphere [56]. These objectives – similarity and uniformity – are quantified by alignment and uniformity scores, respectively. Following Goel et al. [19] and the notations in §3, we compute the alignment score and uniformity score as:
where is the number of text-image pairs.
captures the average cosine similarity between corresponding text and image embeddings, while measures how evenly embeddings are spread across the space. High alignment indicates strong correlation between paired embeddings, whereas low uniformity suggests diverse and efficient embedding distribution—desirable for tasks like retrieval. As shown in Table 5, CLIP achieves stronger alignment but suffers from poor uniformity, leading to redundant representations. On the other hand, RankCLIP along with two of its ablated version, and presents much better balance between alignment and uniformity. These results further suggest that optimizing solely for alignment or uniformity does not necessarily translate to better task performance.
CIFAR-10 | CIFAR-100 | ImageNet1K | ||||
CLIP | 0.28 | -0.19 | 0.28 | -0.18 | 0.33 | -0.19 |
RankCLIP | 0.23 | -0.14 | 0.26 | -0.13 | 0.29 | -0.13 |
0.23 | -0.13 | 0.24 | -0.12 | 0.29 | -0.14 | |
0.25 | -0.15 | 0.28 | -0.14 | 0.32 | -0.17 |
6.3 Qualitative Examples
Class activation maps.
To further examine the effects of ranking consistency, we visualize class activation maps (CAMs) [47] for RankCLIP and CLIP in Fig. 7. The results show that RankCLIP consistently attends to more semantically relevant regions in the images. For example, when given the caption ‘An airport filled with planes sitting on tarmacs’, CLIP mistakenly highlights surrounding areas, whereas RankCLIP focuses precisely on the planes. Similar improvements are observed across other examples, demonstrating that RankCLIP better aligns textual descriptions with visual concepts. This suggests that ranking consistency enhances fine-grained feature learning, leading to more localized and accurate visual grounding.
Text-to-image retrieval.
Fig. 5 compares text-to-image retrieval results from RankCLIP and CLIP. Given a query, we display the top-ranked images retrieved by each model. RankCLIP consistently retrieves more semantically aligned images, even beyond the correct reference image. For instance, in the example of ’A cute cat laying down in a sink’, correctly identifies a cat in a sink, whereas CLIP misidentifies it due to the visual similarity between sinks and toilets. This demonstrates RankCLIP ’s ability to capture fine-grained semantic distinctions, reinforcing its advantage in retrieval tasks that demand precise understanding.
7 Conclusion
In this paper, we introduce RankCLIP, a novel language-image pretraining method that integrates ranking consistency into the contrastive learning paradigm. RankCLIP aims to better understand the complex many-to-many relationships in diverse text-image pairs by optimizing a self-supervised, list-wise rank loss. Through extensive experiments, including zero-shot classification, robustness to distribution shifts, linear probing, and zero-shot image-text retrieval, RankCLIP not only enhances performance but also improves model robustness and semantic comprehension, outperforming the baseline CLIP and another state-of-the-art model ALIP. Our ablation studies and analyses further demonstrate and interpret the significance of each component of RankCLIP in boosting performance and understanding across modalities. We believe that the methodologies and principles of RankCLIP will inspire further research and lead to the development of models with a deeper understanding of the intricate interactions between visual and textual data.
Appendix
Appendix A Training Procedures
A.1 Implementation Details
For CLIP [43], we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP.. And for ALIP [58], we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP.. As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance.. We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted. All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights. And the embedding sizes for both modalities all project to 1024 across the three models.
A.2 Training Parameters
Following CLIP [43], we adopt the ResNet-50 [21] and transformer architectures [13] for image and text encoding, respectively. Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.
A.3 Training Time Consumption
we conducted the experiments using the same hardware specifications. The table below shows the time consumption for training our RankCLIP and CLIP models with 50K samples from CC3M using a single NVIDIA A100 GPU.
Time consumption | Dataset size | epochs | batch_size | model_name | |
CLIP | 1d 2h 54m 48s | 50K | 64 | 512 | RN50 |
RankCLIP | 1d 1h 4m 23s | 50K | 64 | 512 | RN50 |
As shown in the table, the difference in time consumption is negligible. Interestingly, our method is slightly faster than CLIP, but we think it may be attributed to hardware optimizations or variance.
Appendix B CLIP Preliminaries
CLIP [43] has been a prominent method for learning detailed multimodal representations through the alignment of images and texts. Given a set of image-text pairs, where denotes an image and is the corresponding text, the goal is to learn representations that map semantically similar images and texts closer in the embedding space, while dissimilar pairs are distanced apart. More specifically, the foundational CLIP model employs two encoders: an image encoder that processes raw images into visual embeddings and a text encoder which encodes textual data into text embeddings. Then both the text and visual features are projected to a latent space with identical dimension. Formally, the embeddings for a text-image pair are denoted as and , respectively. The embeddings are then normalized to lie on an unit hypersphere by enforcing -norm constraint:
so that the magnitude information is erased and only direction is preserved.
To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss [39], which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
where the first term contrasts images with the texts, the second term contrasts texts with the images, and denotes a temperature scaling parameter that adjusts the concentration of the distribution. The optimization of Eqn. (12) results in embeddings where the cosine similarity between matched image-text pairs is maximized in comparison to unmatched pairs, thus achieving the desired alignment in the joint embedding space.
Despite the efficacy of CLIP in learning correlated multimodal embeddings, it inherently relies on strict pairwise matched comparisons and fails to capture the more complex, fine-grained nature of semantic similarity within and across modalities that are generally treated as unmatched. This observation motivates the development of RankCLIP, which innovates beyond binary pairwise contrasts to consider holistic listwise consistency within and across modalities.
B.1 Additional Experiments
We conduct the linear probing experiment under different training datasize from 3m to 15m as shown in 7.
Data Size | Method | Model Type |
CIFAR-10 |
CIFAR-100 |
DTD |
FGVGAircraft |
Food101 |
GTSRB |
OxfordPets |
SST2 |
STL10 |
SVHN |
CLIP | RN50 | 80.12% | 58.50% | 57.18% | 39.75% | 59.14% | 72.41% | 61.73% | 54.48% | 86.01% | 58.92% | |
3m | RankCLIP | RN50 | 78.29% | 56.24% | 57.82% | 39.30% | 58.63% | 74.13% | 64.35% | 55.02% | 86.69% | 60.68% |
CLIP | ViT-B/32 | 77.60% | 56.15% | 43.19% | 22.59% | 39.72% | 62.05% | 40.39% | 50.96% | 78.99% | 50.53% | |
RankCLIP | ViT-B/32 | 78.42% | 56.64% | 42.39% | 23.43% | 40.19% | 60.63% | 40.56% | 53.32% | 79.60% | 47.72% | |
CLIP | RN50 | 78.81% | 56.32% | 61.49% | 25.83% | 61.64% | 68.76% | 60.37% | 55.57% | 89.82% | 47.99% | |
15m | RankCLIP | RN50 | 83.27% | 62.96% | 65.96% | 32.19% | 68.11% | 74.25% | 67.40% | 56.34% | 94.20% | 53.06% |
CLIP | ViT-B/32 | 82.97% | 62.55% | 49.47% | 24.48% | 52.46% | 63.55% | 50.78% | 52.66% | 87.14% | 46.38% | |
RankCLIP | ViT-B/32 | 82.79% | 59.89% | 52.50% | 23.94% | 56.44% | 61.58% | 52.98% | 53.60% | 89.01% | 42.16% |